OpenStudy (anonymous):

can i write tan(^-1)x, (read as tan inverse x) as 1/tanx ?

OpenStudy (anonymous):

maybe?

zepdrix (zepdrix):

\[\Large \tan^{-1}x \quad\ne\quad \frac{1}{\tan x}\] It's not an exponent when it's -1, even though it's in the same location. I usually try to avoid that notation by using:\[\Large \arctan x\]

OpenStudy (anonymous):

so how would you integrate \[\tan ^{-1}x\] ?

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \arctan x\;dx\]Ummm we'd have to do integration by parts.\[\Large\bf\sf u=\arctan x, \qquad\qquad dv=dx\]

zepdrix (zepdrix):

\[\Large\bf\sf du=?, \qquad\qquad v=?\]

zepdrix (zepdrix):

Have you learned about `Integration by Parts` ? :)

OpenStudy (anonymous):

uh, yes i have.. do you mean i should write \[\tan ^{-1}x . 1\] and then integrate it by parts?

OpenStudy (anonymous):

or there some other method?

zepdrix (zepdrix):

times 1, is that what that notation is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i could take u = tan^-1x and v = 1 and then integrate it... but would that work?

OpenStudy (anonymous):

or will i get stuck somewhere?

zepdrix (zepdrix):

Ok so I guess you learned parts the sloppy way hehe: without using the differential dx. That's fine, we can work with that. So our parts will be: \[\Large\bf\sf u=\tan^{-1}x,\qquad\qquad v'=1\]So then we need to find our:\[\Large\bf\sf u'=?,\qquad\qquad v=?\]In order to do the whole setup, right?

zepdrix (zepdrix):

You remember the derivative of tan^(-1)x? The integral you end up with after doing parts will work out really nicely. It'll be a simple U-substitution.

OpenStudy (anonymous):

so is the integration like this \[xtan ^{-1}x - \frac{ 1}{ 2 }\log_{e}|1+x ^{2}| + c \] ?

zepdrix (zepdrix):

Mmm yes that looks right! good job.

OpenStudy (anonymous):

oh dude!!! this means i didnt get this right in my test this morning. i wrote tan inverse as 1/tan which is cot and then integrated it because it seemed so simple!!... anyway, thanks for your help :)

zepdrix (zepdrix):

oh boy :3